wiki:UC20200407

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 04/07/2020

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), Misty Bentz (GSU), Moire Prescott (NMSU), Mukremin Kilic (OU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Chip Kobulnicky (UWy), Joanne Hughes (Seattle U), Sarah Tuttle (UW), Ben Williams (UW), Derek Buzasi (FGCU)


User feedback and comments from institutional representatives:

  • Colorado - no report
  • FGCU - nothing to report
  • Georgia State - nothing to report
  • JHU - Question from user: Is ARCSAT available if the 3.5m is offline? Answer: No it is not, since the 3.5m observing specialists are needed in order to oversee the operations of ARCSAT and troubleshoot hardware in the dome if necessary, and they are not on-site during the shutdown.
  • NAPG - no report
  • NMSU - nothing to report
  • Oklahoma - Their scheduled student visit didn’t happen due to the COVID-19 shutdown but they conducted two remote training sessions with Candace; they went very well and she did a great job. The students practiced using TUI, they manipulated the instruments and acquired some calibration data. There were 15 students logged in from all different locations. The main challenge was working with that many students at once; for a complete training session Muk would want to limit it to smaller groups of ~5 students. The students couldn’t get on the sky so their feedback was limited, but Muk will ask the students for a more official evaluation of the training process and pass it on the Candace and Nancy. We are eager to explore ways in which we can improve and enhance our remote training offerings, so this feedback will be very useful. We are also checking with Adam at CU to determine whether there were specific aspects of the remote training his students did that were particularly challenging to do remotely; for those we will explore making training videos.
  • Seattle - Joanne got some good data just prior to the shutdown. She noticed a few extra bad columns on ARCTIC and will look at that more closely and provide more details. She also noticed an improvement with the new NA2 baffle in that she was able to get better spatial coverage towards the edges of the field. She will compare this to data taken before the baffle was installed to try and quantify the improvement.
  • UVa - Mike Skrutskie asked if it would be permissible for students in his class to log into TUI and use Tspec as a demo while we are shutdown? Bill replied YES, this is fine.
  • Washington - nothing to report
  • Wyoming - nothing to report

APO and the COVID-19 virus

We are following the guidelines put in place by the NM Governor. The first set of guidelines enabled us to operate with a reduced staff for about a week; subsequent restrictions were put into place that made any work at APO impossible. As of yesterday the Governor extended those restrictions through the end of April. We should be able to stay in this mode indefinitely provided we can still get LN2 to keep the instruments cold. One person visits the site per day to check on things and fill the instruments. To be clear, none of the three ARC telescopes (3.5m, ARCSAT, SDSS) is operating at the moment. Once we get permission to return to the site, we can probably resume 3.5m operations relatively quickly... maybe with as little as one day plus one night of engineering to check everything out. Assuming everything comes back gracefully we will return to science right away with every instrument except for NICFPS, which was warmed up before the shutdown because it requires twice daily LN2 fills, which is not possible given our current situation. While it is warm we will do the servicing of NICFPS that would normally be done over summer shutdown prior to cooling it back down. Plans and schedules for the summer shutdown are still a huge unknown. Mark hasn't reached out to KPNO yet because everything is still so uncertain. We are nominally scheduled to have the primary mirror of the 3.5m realuminized in July, but we are concerned that they may withdraw that option for us once they return to operations and try to get caught up with everything in their queue. Everything is still TBD at this stage.


Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 2/27/2020 - 4/01/2020

0) Overview

The COVID-19 pandemic hit New Mexico and the Governor started putting business restrictions in place by mid-March. For the first week we were able to reduce daytime staffing and keep nighttime observing going. Weather didn’t allow for too many observations but all in all it went fine. Then New Mexico changed and closed all non-essential businesses. Since that time APO has been closed for all observations. Thus we started enacting procedures that placed us in a hold mode, including no observing and absolute minimal staff on site for absolute minimal time. Non-essential systems were powered down or shut off, the telescopes were safed and instruments kept cold and operating except for NIC-FPS (see below). We now have one person visiting the site per day to fill instruments and check the site, plus one person making LN2 deliveries weekly. This helps reduce likelihood of catastrophes and provides some basic security, while not increasing the risk of spreading the virus. Note: these operational procedures are observatory-wide and thus include SDSS operations as well. The one person who comes to the site per day is being rotated among several of the engineering staff between the 3.5m and SDSS.

1) Telescope

Nominal operations with minimal problems. Telescope is safed at zenith.

2) Instruments

DIS is powered on and kept cold via cryotiger chillers.

TripleSpec is powered on and cold.

Agile is powered on and ready.

Echelle is powered on and cold.

NIC-FPS required fills twice daily which was not compatible with a single person coming to the site once per day. Fills were stopped and the instrument was powered off.

ARCTIC is powered on and kept cold via a cryotiger chiller.

ARCSAT: is powered off. Prior to the shut down the new computer was installed in the dome and commissioning started with having two computers splitting functions to improve performance and reduce crashes.


Additional telescope and instrument discussion:

There were no anomalous issues during times we were operating the 3.5m prior to shutdown. For ARCSAT, the installation of the second computer was completed prior to the shutdown. For the Shelyak spectrometer some of the integration had started but then it came to a halt. Derek and Bill discovered that a few components were missing and those have been ordered and are on the way (new guide camera, new fiber). Bill may bring the server home so that he can continue to work on the software development. They expect that once we return to operations it will be few weeks to a month before commissioning the instrument on ARCSAT.


Summary of Open House Event

On March 7 we held a very successful open house event at the observatory. For nighttime activities we took reservations for groups of 12 people to tour the 3.5m and look through the eyepiece (we had to stop observing around 8:30 pm due to precipitation but we still offered tours). We also had a large number of daytime activities - spearheaded by Candace - that were run by the NMSU Astronomy graduate students. During the day we also offered tours of the SDSS facilities and the 3.5m. At night we offered talks at the Sunspot Astronomy and Visitors Center while people were waiting for shuttles to take them over to APO. The NMSU staff that runs the Visitors Center at Sunspot provided tremendous support for this event. All in all it was a hugely successful public outreach event. Mark estimates that all told -- between the daytime and nighttime activities -- we probably had about 500 visitors to the site.


Q2 3.5m scheduling:

When we have knowledge of when we will be reopening the telescope we will revisit the schedule (assuming it happens in Q2 at all). We will review who lost time (how much time was lost, how high a priority the observations were, whether they were time sensitive), and all remaining OPEN/DD time will be assigned to help those who lost time during this shutdown. Fortunately we currently have a fair bit of OPEN/DD time remaining at the end of June since we held those dates for the original summer shutdown schedule (which was subsequently changed to mid-July). In the mean time Ben will attempt to make the Q3 allocations (once we have a better sense of when we will reopen) so he can send them out in late April or early May. Mark asked Nancy how she plans to account for this down time in her annual reporting to the ARC Board of Governors at the end of the year. It is not Engineering, so it is more akin to "weather." But when we calculate the closure statistics we should probably simply reduce the total number of nights in the year (by however many nights we are shut down for); otherwise we will have an artificially high closure rate due to "weather."


Q2 ARCSAT scheduling:

A similar strategy as was described for the 3.5m will be used for dealing with the Q2 ARCSAT schedule once we reopen. This may involve asking some people to share weeks when possible. There were no Q2 student training/classes scheduled for a week of observing. Muk was supposed to get 1 night on ARCSAT in April in between his 3.5m observing nights for his class.


NA2 Baffle Project:

There are no more updates so this item will be removed from future agendas. The baffle is installed and seems to be working well; it will be the default baffle in place unless users explicitly ask for the old baffle (and with enough advanced notice to accommodate the switch). Users are asked to provide any before-and-after comparisons that they can generate to help us quantify the improvement afforded by the new baffle.


KOSMOS Update:

Work on KOSMOS is on hold. We have detector back and in lab. We are still seeing some anomalous behavior with the dewar; Rishi and Kal did some testing with a borrowed helium leak detector; all of those data are now accumulated in one place and Sarah will be going through it. UW is on a stay at home order until May 4. Hopefully after that the TEG members can go back to the lab to reassemble KOSMOS in the new dewar; they are making preparations now so they can do that successfully when the time comes.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: No update this month. Nancy will contact Adam and suggest they schedule another meeting.
  • Getting narrow-band filters for ARCTIC: in progress. Nancy is having difficulty placing the order given that the vendor is in Japan and we are in a pandemic.

ACTION ITEMS:

Open action items from previous meetings:

  • Nancy will contact trip leaders bringing student groups to APO in Q1 (-> Q2) to discuss training video ideas. STATUS: OPEN. Nancy will contact Candace and Adam Kowalski from CU, who are planning to conduct a fully online training session, to find out what was missing or most challenging that would have benefited from having some video tutorials. She will communicate that info to Muk (OU), who is bringing a training group to the site in April, so his students can focus on those topics. UPDATE: Candace has contacted Adam about this and is awaiting a reply. Muk provided feedback from his recent trip. We will table this until we receive more explicit feedback about which topics would be appropriate for training videos. STATUS: CLOSED.
  • UC reps should notify their users of the NA2 baffle switch and request feedback on performance, particularly if any users have before-and-after data that they can use for a quantitative comparison. STATUS: CLOSED.
  • UC reps should ask if any of their users would like to help with the commissioning of the Shelyak echelle spectrometer on ARCSAT. Bill has received several names. STATUS: CLOSED.

New action items from this meeting:

  • None.

All Other Business

Nancy is having the Science Communications Interns working with her start designing a new brochure since the current one is fairly outdated. We discussed whether or not to have a single APO-wide brochure and/or a 3.5m brochure. The suggestion was that we keep the brochure APO-wide, and we develop a different document (e.g. a fact sheet) that is solely about the 3.5m.


Next meeting:

The next Users Committee Zoom meeting will be on Monday May 4 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.